# UDC: A Unified Neural Divide-and-Conquer Framework for Large-Scale Combinatorial Optimization Problems

### v1

@@ Abstract
Single-stage neural combinatorial optimization solvers have achieved near-optimal results on various small-scale combinatorial optimization (CO) problems without needing expert knowledge. However, these solvers exhibit significant performance degradation when applied to large-scale CO problems. Recently, two-stage neural methods with divide-and-conquer strategies have shown superiorities in addressing large-scale CO problems. Nevertheless, the efficiency of these methods highly relies on problem-specific heuristics in either the divide or the conquer procedure, which limits their applicability to general CO problems. Moreover, these methods employ separate training schemes and ignore the interdependencies between the dividing and conquering strategies, which often leads to sub-optimal solutions. To tackle these drawbacks, this article develops a unified neural divide-and-conquer framework (i.e., UDC) for solving general large-scale CO problems. UDC offers a Divide-Conquer-Reunion (DCR) training method to eliminate the negative impact of a sub-optimal dividing policy. Employing a high-efficiency Graph Neural Network (GNN) for global dividing and a fixed-length sub-path solver for conquering sub-problems, the proposed UDC framework demonstrates extensive applicability, achieving superior performance in 10 representative large-scale CO problems.
@@ 1 Introduction
Combinatorial optimization (CO) [1] has numerous practical applications including route planning [2], circuit design [3], biology [4], etc. Over the past few years, neural combinatorial optimization (NCO) methods have been developed to efficiently solve typical CO problems such as the Travelling Salesman Problem (TSP) and Capacitated Vehicle Routing Problem (CVRP). Among them, reinforcement learning (RL)-based constructive NCO methods [5, 6] have shown the performance of generating near-optimal solutions for small-scale instances (e.g., TSP instances with no more than 200 nodes) without the requirement of expert knowledge [7]. These solvers mainly adopt a single-stage solution generation process, constructing the solution of CO problems step-by-step in an end-to-end manner. However, hindered by the disability of training on larger-size instances [8, 9, 10], these methods show limited performance when generalizing to solve large-scale CO instances.
@@
To meet the demands of larger-scale CO applications, researchers have paid increasing attention to extending NCO methods to larger-scale instances (i.e., 1,000-node ones) [11]. Existing NCO methods for large-scale problems mainly involve modified single-stage solvers [6] and neural divide-and-conquer methods [12, 13]. BQ-NCO [14] and LEHD [8] develop a sub-path construction process [15] and employ models with heavy-decoder to learn the process. Nevertheless, training such a heavyweight model necessitates supervised learning (SL), which limits their applicability to problems where high-quality solutions are not available as labels. ELG [6], ICAM [16], and DAR [17] integrate auxiliary information to guide the learning of RL-based single-stage constructive solvers [5]. However, the auxiliary information needs problem-specific designs, and the $\mathcal{O}(N^{2})$ complexity caused by the self-attention mechanism also poses challenges for large-scale ($N$-node) problems. As another category of large-scale NCO, neural divide-and-conquer methods have attracted increasing attention. Inspired by the common divide-and-conquer idea in traditional heuristics [18, 19], these methods execute a two-stage solving process: overall instance partition (i.e., dividing stage) and sub-problem solving (i.e., conquering stage). Using neural network learning in one or both stages, TAM [12], H-TSP [20], and GLOP [13] have demonstrated superiority in large-scale TSP and CVRP.
@@
Despite acknowledging the prospect, existing neural divide-and-conquer approaches still suffer from shortcomings in applicability and performance. Some neural divide-and-conquer methods rely on problem-specific heuristics in either dividing (e.g., GLOP [13] and SO [21]) or conquering stages (e.g., L2D [22] and RBG [23]), which limits their generalizability across other CO problems. Moreover, all the existing neural divide-and-conquer methods adopt a separate training process, training the dividing policy after a pre-trained conquering policy [13, 12, 20]. However, the separate training scheme overlooks the interdependencies between the dividing and conquering policies and may lead to unsolvable antagonisms. Once the pre-trained conquering policy is sub-optimal for some sub-problems, such a training scheme might guide the dividing policy to adapt the pre-trained sub-optimal conquering policy, thus converging to local optima.
@@
As shown in Table 1, this paper finds that considering the negative impact of sub-optimal dividing results is essential in generating divide-and-conquer-based solutions. To this aim, we propose a novel RL-based training method termed Divide-Conquer-Reunion (DCR). Enabling the DCR in a unified training scheme, we propose the unified neural divide-and-conquer framework (UDC) for a wide range of large-scale CO problems. UDC employs a fast and lightweight GNN to decompose a large-scale CO instance into several fixed-length sub-problems and utilizes constructive solvers for solving these sub-problems. We conduct extensive experiments on 10 different CO problems. Experimental results indicate that UDC can significantly outperform existing large-scale NCO solvers in both efficiency and applicability without relying on any heuristic design.
@@
The contributions are as follows: 1) We propose a novel method DCR to enhance training by alleviating the negative impact of sub-optimal dividing policies. 2) Leveraging DCR in training, the proposed UDC achieves a unified training scheme with significantly superior performance 3) UDC demonstrates extensive applicability and can extend to general CO problems with similar settings.
@@ 2.1 Divide-and-Conquer
where $\mathcal{G}$ represents the problem instance and ${\Omega}$ is a set consisting of all feasible solutions.
Leveraging the divide-and-conquer idea in solving CO problems is promising in traditional (meta) heuristic algorithms [25] especially large-neighborhood-search-based algorithms [26, 27]. These methods achieve gradual performance improvement from an initial solution through iterative sub-problem selection and sub-problem repair. Recent neural divide-and-conquer methods conduct a similar process and employ efficient deep learning techniques for initial solution, sub-problem partition (i.e., dividing policy), and sub-problem solving (i.e., conquering policy) [13]. The dividing policy $\pi_{D}(\mathcal{G})$ decomposes the instance $\mathcal{G}$ to $K$ mutually unaffected sub-problems $\{\mathcal{G}_{1},\mathcal{G}_{2},\ldots,\mathcal{G}_{K}\}$ and the conquering policy $\pi_{C}$ generates the solutions $\boldsymbol{s}_{k},k\in\{1,\ldots,K\}$ of corresponding sub-problems by $\boldsymbol{s}_{k}\sim\pi_{C}(\mathcal{G}_{k})$. Finally, a total solution $\boldsymbol{x}$ of $\mathcal{G}$ is merged as $\boldsymbol{x}=\text{Concat}(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{K})$.
2.2 Constructive Neural Solvers
The (single-stage) constructive NCO solver has garnered attention from researchers for its ability to solve general CO problems with time efficiency. These constructive solvers typically employ an attention-based encoder-decoder network structure, where a multi-layer encoder processes CO instances into embeddings in hidden spaces and a lightweight decoder handles dynamic constraints while constructing feasible solutions [5]. With being modeled as Markov Decision Processes (MDPs) [28], constructive solvers can be trained using deep reinforcement learning (DRL) techniques without requiring expert experience. In the solution generation process, constructive solvers with parameter $\theta$ process the policy $\pi$ of a $\tau$-length solution $\boldsymbol{x}=\{x_{1},\ldots,x_{\tau}\}$ as follows:
@@ 2.2 Constructive Neural Solvers
where $\boldsymbol{x}_{1:t-1}$ represents the partial solution before the selection of $x_{t}$. DRL-based constructive NCO solvers demonstrate outstanding solution qualities in small-scale CO problems. So existing neural divide-and-conquer methods [13, 21] generally employ a pre-trained constructive neural solver (e.g., Attention Model [28]) for their conquering policy $\pi_{C}$.
2.3 Heatmap-based Solver
Advanced constructive NCO solvers rely on the self-attention encoder layers [29] for node representation, whose quadratic time and space complexity related to problem scales [30] hinders the direct training on large-scale instances. So, heatmap-based solvers [31, 32] are proposed to solve large-scale CO problems efficiently by a lighter-weight GNN [33]. In heatmap-based solvers for VRPs, the GNN with parameter $\phi$ first generates a heatmap $\mathcal{H}\in\mathbb{R}^{N\times N}$ within linear time [34] and a feasible solution is then constructed based on this heatmap with the policy [35] as follows:
@@ 2.3 Heatmap-based Solver
However, the non-autoregressively generated heatmap [36] excludes the representations of the constructed partial solution $\boldsymbol{x}_{1:t-1}$, so the greedy decoding policy in heatmap-based solvers becomes low-quality and these solvers highly rely on search algorithms [37, 38] for higher-quality solutions.
Figure 1: The proposed UDC Framework.
@@ 3 Methodology: Unified Divide-and-Conquer
As shown in Figure 1, UDC follows the general framework of neural divide-and-conquer methods [13], solving CO problems through two distinct stages: the dividing stage and the conquering stage. The dividing stage of UDC is expected to generate a feasible initial solution with a similar overall shape compared to the optimal solution, so the subsequent conquering stage is likely to obtain more optimal solutions by solving local sub-problems. For all the involved CO problems, UDC first constructs a sparse graph to represent the problem instance and then employs a heatmap-based solver with Anisotropic Graph Neural Network (AGNN) [35, 39, 40] to process the sparse graph and generate initial solutions based on the dividing policy. The following conquering stage first decomposes the original problem instance into multiple sub-problems based on the initial solution while integrating necessary constraints. Then, towards minimizing the decomposed objective function, constructive-based solvers are utilized as the conquering policy in generating sub-solutions of these sub-problems. There is no single universal constructive solver that can handle all types of CO problems [41], so we employ suitable constructive neural solvers for each sub-CO-problem (i.e., AGNN [39] for MIS, POMO [5] or ICAM [16] for VRPs).
@@ 3.1 Pipeline: Dividing Stage & Conquering Stage
Conquering Stage: Sub-problem Preparation: The conquering stage first generates pending sub-problems along with their specific constraints. For VRPs, the nodes in sub-VRPs and the constraints of sub-VRPs are built based on continuous sub-sequences in the initial solution $\boldsymbol{x}_{0}$. In generating $n$-node sub-problems, UDC divides the original $N$-node instance $\mathcal{G}$ into $\lfloor\frac{N}{n}\rfloor$ sub-problems $\{\mathcal{G}_{1},\ldots,\mathcal{G}_{\lfloor\frac{N}{n}\rfloor}\}$ according to $\boldsymbol{x}^{0}$, temporarily excluding sub-problems with fewer than $n$ nodes. The constraints $\{\Omega_{1},\ldots,\Omega_{\lfloor\frac{N}{n}\rfloor}\}$ of sub-problems include not only the problem-specific constraints (e.g., no self-loop in sub-TSPs) but also additional constraints to ensure the legality of the merged solution after replacing the original solution sub-sequence with the sub-problem solution (e.g., maintaining the first and last node in sub-VRPs). Similar to previous works [13, 42], we also normalize the coordinate of sub-problems and some numerical constraints to enhance the homogeneity of pending sub-problem instances. Due to the need to handle different constraints, the sub-problem preparation processes vary for different CO problems, as listed in the Appendix B.
Conquering Stage: Constructive Neural Conquering: After the preparation of sub-problems, the conquering policy is utilized to generate solutions to these instances $\{\mathcal{G}_{1},\ldots,\mathcal{G}_{\lfloor\frac{N}{n}\rfloor}\}$. We utilize a constructive solver with parameter $\theta$ for most involved sub-CO problems. Their sub-solution $\boldsymbol{s}_{k}=(s_{k,1},\ldots,s_{k,n}),k\in\{1,\ldots,\lfloor\frac{N}{n}\rfloor\}$ are sampled from the conquering policy $\pi_{c}$ as follows:
$\pi_{c}(\boldsymbol{s}_{k}|\mathcal{G}_{k},\Omega_{k},\theta)$ =$&\prod_{t=2}^{n}p(\boldsymbol{s}_{k,t}|\boldsymbol{s}_{k,1:t-1}\mathcal{G}_{k},\Omega_{k},\theta),&\text{if}\ \boldsymbol{s}_{k}\in\Omega_{i}\\ &0,&\text{otherwise}\end{aligned}\right.,$ (5)
where $\boldsymbol{s}_{k,1:t-1}$ represents the partial solution before the selection of $s_{k,t}$. Finally, as the last step in a conquering stage, the sub-solution with improvement on the objective function will replace the original solution fragment in $\boldsymbol{x}_{0}$ and the merged solution becomes $\boldsymbol{x}_{1}$. The conquering stage can execute repeatedly on the new merged solution for a gradually better solution quality and we note the solution after $r$ conquering stages as $\boldsymbol{x}_{r}$.
Figure 2: The proposed DCR training method.

### v2

@@ Abstract
Single-stage neural combinatorial optimization solvers have achieved near-optimal results on various small-scale combinatorial optimization (CO) problems without needing expert knowledge. However, these solvers exhibit significant performance degradation when applied to large-scale CO problems. Recently, two-stage neural methods with divide-and-conquer strategies have shown efficiency in addressing large-scale CO problems. Nevertheless, the performance of these methods highly relies on problem-specific heuristics in either the divide or the conquer procedure, which limits their applicability to general CO problems. Moreover, these methods employ separate training schemes and ignore the interdependencies between the dividing and conquering strategies, which often leads to sub-optimal solutions. To tackle these drawbacks, this article develops a unified neural divide-and-conquer framework (i.e., UDC) for solving general large-scale CO problems. UDC offers a Divide-Conquer-Reunion (DCR) training method to eliminate the negative impact of a sub-optimal dividing policy. Employing a high-efficiency Graph Neural Network (GNN) for global instance dividing and a fixed-length sub-path solver for conquering divided sub-problems, the proposed UDC framework demonstrates extensive applicability, achieving superior performance in 10 representative large-scale CO problems. The code is available at https://github.com/CIAM-Group/NCO_code/tree/main/single_objective/UDC-Large-scale-CO-master
@@ 1 Introduction
Combinatorial optimization (CO) [1] has numerous practical applications including route planning [2], circuit design [3], biology [4], etc. Over the past few years, neural combinatorial optimization (NCO) methods are developed to efficiently solve typical CO problems such as the Travelling Salesman Problem (TSP) and Capacitated Vehicle Routing Problem (CVRP). Among them, reinforcement learning (RL)-based constructive NCO methods [5, 6] can generate near-optimal solutions for small-scale instances (e.g., TSP instances with no more than 200 nodes) without the requirement of expert knowledge [7]. These solvers usually adopt a single-stage solution generation process, constructing the solution of CO problems node-by-node in an end-to-end manner. However, hindered by the disability of training on larger-size instances directly [8, 9, 10], these methods exhibit limited performance when generalizing to solve large-scale CO problems.
@@
To meet the demands of larger-scale CO applications, researchers have increasingly focused on extending NCO methods to larger-scale instances (e.g., 1,000-node ones) [11]. Existing NCO methods for large-scale problems primarily involve modified single-stage solvers [6] and neural divide-and-conquer methods [12, 13]. BQ-NCO [14] and LEHD [8] develop a sub-path construction process [15] and employ models with heavy-decoder to learn the process. Nevertheless, training such a heavyweight model necessitates supervised learning (SL), which limits their applicability to problems where high-quality solutions are not available as labels. ELG [6], ICAM [16], and DAR [17] integrate auxiliary information to guide the learning of RL-based single-stage constructive solvers [5]. However, the auxiliary information needs problem-specific designs, and the $\mathcal{O}(N^{2})$ complexity caused by the self-attention mechanism also poses challenges for large-scale ($N$-node) problems. As another category of large-scale NCO, neural divide-and-conquer methods have attracted increasing attention. Inspired by the common divide-and-conquer idea in traditional heuristics [18, 19], these methods execute a two-stage solving process, including a dividing stage for overall instance partition and a conquering stage for sub-problem solving. By employing neural network for one or both stages, TAM [12], H-TSP [20], and GLOP [13] have demonstrated superior efficiency in large-scale TSP and CVRP.
@@
Despite acknowledging the prospect, these neural divide-and-conquer approaches still suffer from shortcomings in applicability and solution quality. Some neural divide-and-conquer methods rely on problem-specific heuristics in either dividing (e.g., GLOP [13] and SO [21]) or conquering stages (e.g., L2D [22] and RBG [23]), which limits their generalizability across various CO problems. Moreover, all the existing neural divide-and-conquer methods adopt a separate training process, training the dividing policy after a pre-trained conquering policy [13, 12, 20]. However, the separate training scheme overlooks the interdependencies between the dividing and conquering policies and may lead to unsolvable antagonisms. Once the pre-trained conquering policy is sub-optimal for some sub-problems, such a training scheme might guide the dividing policy to adapt the pre-trained sub-optimal conquering policy, thus converging to local optima.
@@
As shown in Table 1, this paper finds that considering the negative impact of sub-optimal dividing results is essential in generating high-quality divide-and-conquer-based solutions. To this aim, we propose a novel RL-based training method termed Divide-Conquer-Reunion (DCR). Enabling the DCR in a unified training scheme, we propose the unified neural divide-and-conquer framework (UDC) for a wide range of large-scale CO problems. UDC employs a fast and lightweight GNN to decompose a large-scale CO instance into several fixed-length sub-problems and utilizes constructive solvers to solve these sub-problems. We conduct extensive experiments on 10 different CO problems. Experimental results indicate that UDC can significantly outperform existing large-scale NCO solvers in both efficiency and applicability without relying on any heuristic design.
@@
The contributions of this paper are as follows: 1) We propose a novel method DCR to enhance training by alleviating the negative impact of sub-optimal dividing policies. 2) Leveraging DCR in training, the proposed UDC achieves a unified training scheme with significantly superior performance. 3) UDC demonstrates extensive applicability and can be applied to general CO problems with similar settings.
@@ 2.1 Divide-and-Conquer
where $\mathcal{G}$ represents the CO instance and ${\Omega}$ is a set consisting of all feasible solutions.
Leveraging the divide-and-conquer idea in solving CO problems is promising in traditional (meta)heuristic algorithms [25], especially large-neighborhood-search-based algorithms [26, 27]. These methods achieve gradual performance improvement from an initial solution through iterative sub-problem selection (i.e., the dividing stage) and sub-problem repair (i.e., the conquering stage). Recent neural divide-and-conquer methods conduct a similar solving process and employ efficient deep-learning techniques to model the policies of both stages [13]. The dividing policy $\pi_{d}(\mathcal{G})$ decomposes the instance $\mathcal{G}$ to $K$ mutually unaffected sub-problems $\{\mathcal{G}_{1},\mathcal{G}_{2},\ldots,\mathcal{G}_{K}\}$ and the conquering policy $\pi_{c}$ generates sub-solutions $\boldsymbol{s}_{k},k\in\{1,\ldots,K\}$ by $\boldsymbol{s}_{k}\sim\pi_{c}(\mathcal{G}_{k})$ to minimize the objective function $f^{\prime}(\boldsymbol{s}_{k},\mathcal{G}_{k})$ of corresponding sub-problems. Finally, a total solution $\boldsymbol{x}$ of $\mathcal{G}$ is merged as $\boldsymbol{x}=\text{Concat}(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{K})$.
2.2 Constructive Neural Solver
The (single-stage) constructive NCO solvers are promising in solving general small-scale CO problems with time efficiency. These constructive solvers typically employ an attention-based encoder-decoder network structure, where a multi-layer encoder processes CO instances into embeddings in hidden spaces and a lightweight decoder handles dynamic constraints while constructing feasible solutions [5]. With the solving process modeled as Markov Decision Processes (MDPs) [28], constructive solvers can be trained using deep reinforcement learning (DRL) techniques without requiring expert experience. In the solution generation process, constructive solvers with parameter $\theta$ process the policy $\pi$ of a $\tau$-length solution $\boldsymbol{x}=\{x_{1},\ldots,x_{\tau}\}$ as follows:
@@ 2.2 Constructive Neural Solver
where $\boldsymbol{x}_{1:t-1}$ represents the partial solution before the selection of $x_{t}$. DRL-based constructive NCO solvers demonstrate outstanding solution qualities in small-scale CO problems. So existing neural divide-and-conquer methods [13, 21] also generally employ a pre-trained constructive neural solver (e.g., Attention Model [28]) for their conquering policy $\pi_{c}$.
2.3 Heatmap-based Neural Solver
Advanced constructive NCO solvers rely on the self-attention layers [29] for node representation, whose quadratic time and space complexity related to problem scales [30] hinders the direct training on large-scale instances. So, heatmap-based solvers [31, 32] are proposed to solve large-scale CO problems efficiently by a lighter-weight GNN [33]. In heatmap-based solvers for vehicle routing problems (VRPs), a GNN with parameter $\phi$ first generates a heatmap $\mathcal{H}\in\mathbb{R}^{N\times N}$ within linear time [34] and a feasible solution is then constructed based on this heatmap with the policy [35] as follows:
@@ 2.3 Heatmap-based Neural Solver
However, the non-autoregressively generated heatmap [36] excludes the information about the order of the constructed partial solution $\boldsymbol{x}_{1:t-1}$, so the greedy decoding policy in heatmap-based solvers becomes low-quality and these solvers rely on search algorithms [37, 38] for higher-quality solutions.
Figure 1: The solving process of the proposed UDC on large-scale VRP, 0-1 Knapsack Problem (KP), and Maximum Independent Set (MIS).
@@ 3 Methodology: Unified Divide-and-Conquer
As shown in Figure 1, UDC follows the general framework of neural divide-and-conquer methods [13], solving CO problems through two distinct stages: the dividing stage and the conquering stage. The dividing stage of UDC is expected to generate a feasible initial solution with a similar overall shape compared to the optimal solution, so the subsequent conquering stage can obtain more high-quality solutions by solving local sub-problems. For all the involved CO problems, UDC first constructs a sparse graph to represent the instance and then employs a heatmap-based solver with Anisotropic Graph Neural Network (AGNN) as the dividing policy [39, 40] to generate initial solutions. The following conquering stage first decomposes the original CO instance into multiple sub-problems based on the initial solution while integrating the necessary constraints for sub-problems. Then, towards minimizing the decomposed objective function of the sub-problem, UDC uses constructive neural solvers as the conquering policy to generate sub-solutions. There is no single universal constructive neural solver that can handle all types of CO problems [41], so we employ suitable solvers for each sub-CO-problem (e.g., AGNN [39] for MIS, POMO [5] or ICAM [16] for VRPs).
@@ 3.1 Pipeline: Dividing Stage & Conquering Stage
Conquering Stage: Sub-problem Preparation: The conquering stage first generates pending sub-problems along with their specific constraints. For VRPs, the nodes in sub-VRPs and the constraints of sub-VRPs are built based on continuous sub-sequences in the initial solution $\boldsymbol{x}_{0}$. In generating $n$-node sub-problems, UDC divides the original $N$-node instance $\mathcal{G}$ into $\lfloor\frac{N}{n}\rfloor$ sub-problems $\{\mathcal{G}_{1},\ldots,\mathcal{G}_{\lfloor\frac{N}{n}\rfloor}\}$ according to $\boldsymbol{x}^{0}$, temporarily excluding sub-problems with fewer than $n$ nodes. The constraints $\{\Omega_{1},\ldots,\Omega_{\lfloor\frac{N}{n}\rfloor}\}$ of sub-problems include not only the problem-specific constraints (e.g., no self-loop in sub-TSPs) but also additional constraints to ensure the legality of the merged solution after integrating the sub-problem solution to the original solution (e.g., maintaining the first and last node in sub-VRPs). Similar to previous works [13, 42], we also normalize the coordinate of sub-problems and some numerical constraints to enhance the homogeneity of pending sub-problem instances. Due to the need to handle different constraints, the sub-problem preparation processes vary for different CO problems, which is presented in detail in the Appendix B.
Conquering Stage: Constructive Neural Conquering: After the preparation of sub-problems, the conquering policy is utilized to generate solutions to these instances $\{\mathcal{G}_{1},\ldots,\mathcal{G}_{\lfloor\frac{N}{n}\rfloor}\}$. We utilize constructive solvers with parameter $\theta$ for most involved sub-CO problems. Their sub-solution $\boldsymbol{s}_{k}=(s_{k,1},\ldots,s_{k,n}),k\in\{1,\ldots,\lfloor\frac{N}{n}\rfloor\}$ are sampled from the conquering policy $\pi_{c}$ as follows:
$\pi_{c}(\boldsymbol{s}_{k}|\mathcal{G}_{k},\Omega_{k},\theta)$ =$&\prod_{t=1}^{n}p(\boldsymbol{s}_{k,t}|\boldsymbol{s}_{k,1:t-1}\mathcal{G}_{k},\Omega_{k},\theta),&\text{if}\ \boldsymbol{s}_{k}\in\Omega_{k}\\ &0,&\text{otherwise}\end{aligned}\right.,$ (5)
where $\boldsymbol{s}_{k,1:t-1}$ represents the partial solution before the selection of $s_{k,t}$. Finally, as the last step in a conquering stage, sub-solutions with improvements on the objective function will replace the original solution fragment in $\boldsymbol{x}_{0}$ and the merged solution becomes $\boldsymbol{x}_{1}$. It’s worth noting that the conquering stage can execute repeatedly on the new merged solution for a gradually better solution quality and we note the solution after $r$ conquering stages as $\boldsymbol{x}_{r}$.
Figure 2: The proposed DCR training method. The upper part shows the pipeline of a single DCR-enabled training step, including a Divide step, a Conquer step, and a Reunion step. The lower part provides a local view of a solution fragment to demonstrate our motivation, the additional Reunion step can repair the wrong connections from sub-optimal sub-problem dividing results. DCR has the potential to correct sub-optimal dividing results by repairing the connection part in the additional Reunion step.